We are evaluating a project that costs $2,070,000, has a 7-year life, and has no salvage value. Assume that depreciation is straight-line to zero over the life of the project. Sales are projected at 93,400 units per year. Price per unit is $38.73, variable cost per unit is $23.85, and fixed costs are $854,000 per year. The tax rate is 23 percent and we require a return of 12 percent on this project. Suppose the projections given for price, quantity, variable costs, and fixed costs are all accurate to within ±10 percent. Calculate the best-case and worst-case NPV figures.
Answer #1

Best case NPV:
=-2070000+((93400*1.1*(38.73*1.1-23.85*0.9)-854000*0.9-2070000/7)*(1-23%)+2070000/7)/12%*(1-1/1.12^7)=3171087.55226964

Worst case NPV:
=-2070000+((93400*0.9*(38.73*0.9-23.85*1.1)-854000*1.1-2070000/7)*(1-23%)+2070000/7)/12%*(1-1/1.12^7)=-2513845.48555888